# A Mess of Memory System Benchmarking, Simulation and Application Profiling

### v1

@@ II-D Performance characterization: Actual systems
All platforms except AMD Zen2 CPU and NVIDIA H100 show a similar saturated bandwidth range, between approximately 70% and 90% of the maximum theoretical bandwidth. The maximum achieved bandwidth cannot reach the theoretical one because part of it is “lost” due to factors such as DRAM refresh cycles blocking the entire chip, page misses causing precharge and activate cycles, and timing restrictions at the bank, rank, and channel levels [64]. The efficiency of the memory bandwidth utilization also depends on the memory controller design. As expected (see Section II-C), the best utilization is achieved for 100%-read memory traffic, and it reduces with the increment of the memory writes. AMD Zen2 is an exception in two ways. First, its saturated bandwidth range is significantly lower, 57–71% of the maximum theoretical one. Second, it does not follow the expected impact of the write traffic on the bandwidth utilization. Instead, the traffic with the maximum rate of memory writes shows a very good performance, very close to the 100%-read traffic, while the main drop is detected for a mixed, e.g. 60%-read/40%-write, traffic.
We detect the memory system over-saturation, in some AMD Zen 2, Intel Skylake and Intel Cascade Lake bandwidth–latency curves. In Amazon Graviton 3, Intel Sapphire Rapids and NVIDIA H100 the over-saturation is frequent for the memory traffic with high percent of writes. [^2] These findings are consistent with two recent studies which report that running high-bandwidth benchmarks on all CPU cores may lead to lower memory bandwidths w.r.t. the experiment in which some cores are not used [6, 65]. We explored this behavior in the Cascade Lake servers in which we had access to additional memory system counters, e.g. monitoring the row-buffer statistics. In all the memory over-saturation experiments we detect a significant increase in row-buffer misses. In case of a row-buffer miss, the current content of the row is stored in the memory array and the correct row is loaded into the row-buffer. These additional operations increase the memory access time and reduce the effective device bandwidth, matching the over-saturated memory behavior. To confirm these findings, we increased the memory system pressure by removing four out of six DIMMs in each socket. In these experiments, we detected a large over-saturation areas in all bandwidth–latency curves. And indeed, the measurement confirmed that the over-saturation is highly correlated with the row-buffer miss rate.
@@
Finally, we detect that in most of the platforms the bandwidths reported by the STREAM kernels are significantly below the Mess measurements. [^3] The STREAM benchmark measures the performance of four kernels: Copy, Scale, Add and Triad. All the results presented in this study follow the Stream guidelines: they are obtained with the C version of the benchmark running with 64-bit data types with no significantly modified code or assembly language [66]., [^4] In Graviton 3 and NVIDIA H100, the results reported by STREAM are very close to the maximum Mess measurements for the corresponding read/write ratio. This would correspond to a architecture with a write-through cache policy. We did not find any public document that details this aspect of the Graviton 3 and NVIDIA H100 cache organization. In the Intel and A64FX platforms, the STREAM reports around half of the maximum theoretical bandwidth. In IBM Power 9, the STREAM reports only one third of it. There are two main sources of this gap. The first one is that the STREAM does not measure, but estimates the memory bandwidth. This estimate is based on the application execution time, size of the data structures used in the computation, and number of load and store instruction in each STREAM kernel. In particular, STREAM calculation assumes one memory read for each load instruction and one memory write for each store. This is not the case in most of the state-of-the-art HPC servers because they deploy a write-allocate cache policy. Each store instruction, therefore, does not correspond to a single memory write, as assumed in the STREAM bandwidth calculation, but to one read and one write (see Section II-A). The second cause of the gap between the Mess and STREAM bandwidth is the composition of the memory traffic. The Mess benchmark achieves the maximum bandwidth for a 100%-read traffic. The write memory traffic, present in all STREAM kernels, adds timing constrains and reaches sooner the bandwidth saturation point.
II-E Performance characterization: Memory simulators
@@
Mess benchmark can also be used to characterize memory simulators and compare them with the actual systems they intent to model. We illustrate this Mess capability with the gem5, ZSim, and OpenPiton Metro-MPI simulators with different internal memory models and widely-used external memory simulators, DRAMsim3 and Ramulator.
II-E1 gem5
@@
The gem5 [16] is a cycle-accurate full-system simulator. In our experiments, the simulator is configured to model the Graviton 3 server with 64 Neoverse N1 cores [59]. The cache hierarchy includes 64 KB of 4-way L1 instruction and data cache, 1 MB of 8-way private L2 cache and 64 MB of 16-way shared L3. The main memory system has eight eight DDR5-4800 memory channels. Figure 4 compares Mess bandwidth–latency curves of the actual server with gem5 internal memory models: simple memory model and internal DDR model. To maintain a reasonable simulation time, we model each system with a family of six curves, from 50% to 100% read memory traffic with a 10% step.
@@
Practically in the whole bandwidth range, the gem5 simple memory model delivers a fixed latency of 4–49 ns. The latency increases only when the bandwidth asymptotically approaches its theoretical maximum. Contrary to the Graviton 3 server, the highest latencies are measured for a 100%-read traffic, and the latency drops with the percent of memory writes. Also, unlike in the actual platform, for some memory traffic increasing the bandwidth reduces the memory access latency. For example, the 50%-read/50%-write traffic reaches the lowest simulated latency of only 4 ns at the 200 GB/s bandwidth. The same traffic in the actual system has the memory access latency of 261 ns.
@@
The more detailed internal DDR model shows small improvements over the simple memory model, but still poorly resembles the actual system behavior. The simulated latencies are unrealistically low, most of them in the range of 14–100 ns. Similarly to the gem5 simple memory model, the latencies drops with the percent of memory writes. For all the curves exept 100%-read, the saturated bandwidth is significantly lower from the one measured on the actual system. Again, the error increases with the percent of memory writes.
@@
Figure 4: The gem5 memory models poorly resemble the actual system performance. The simulated memory access latency is unrealistically low, starting at 4 ns. Contrary to the actual system, the latencies drop with the percent of memory writes.
@@
Figure 5: Memory system performance of the actual Intel Skylake server and various ZSim memory models.
II-E2 ZSim
We select ZSim [15] as a representative of event-based hardware simulators. We use publicly-available ZSim modeling 24-core Intel Skylake processor connected to six DDR4-2666 channels [67]. The cache hierarchy of the modeled CPU includes 64 KB of 8-way L1 instruction and data cache, 1 MB of 16-way private L2 cache and 33 MB of 11-way shared L3. The simulator is extensively evaluated against the actual hardware platform [68]. The ZSim comprises three internal memory models: fixed-latency, M/D/1 queue model and the internal DDR model. Also it is already connected to Ramulator [24, 69] and DRAMsim3 [22]. To avoid any error due to the integration of an event-based CPU simulator with cycle-accurate memory models [46, 70, 71], we use the ZSim interfaces included in the Ramulator and DRAMsim3 releases.
We compare the Mess bandwidth–latency curves of the actual server with all five ZSim approaches for the main memory simulation (see Figure 5). We configure the fixed-latency model with the unloaded memory latency measured in the actual system. The M/D/1 queue model also requires user to specify the maximum theoretical bandwidth, which is 128 GB/s in the platform under study. The internal DDR model, Ramulator and DRAMsim3 are based on the detailed simulation of the DDR4-2666.
As expected, the fixed-latency memory model provides a constant latency in the whole bandwidth domain. Given that this latency is configured by a user, it can be set to match the unloaded memory latency in the actual system. On the down side, the memory bandwidth provided by this model is unrealistic: the maximum simulated bandwidth is 342 GB/s, which exceeds the maximum theoretical one by 2.7$\times$. The M/D/1 queues correctly model the memory system behavior in the linear part of the curves. The modeling of the system saturation is less accurate. The queue model does show some difference between read and write memory traffic, but the reported performance does not correspond to the actual system trend in which increasing the write traffic lowers the performance. The internal DDR model shows the closest performance match to the actual system — it models the linear and saturated segments of the curves, and the impact of the memory writes. Still, there are three main difference w.r.t. to the actual system. First, the simulator underestimates the saturated bandwidth area to 69–93 GB/s, significantly below the 92–116 GB/s measured in the actual system. Second, the simulator excessively penalizes the memory writes which is seen as a wider spread of the curves with a higher write memory traffic. Third, we detect some unrealistic memory-latency peaks in the low-bandwidth 1–4 GB/s curve segments. The DRAMsim3 shows a similar trend as the M/D/1 queue model in the linear segments of the memory curves, with some latency error, 52–63 ns in the DRAMsim3 versus 89–109 ns in the actual system. The simulator does not model the saturated bandwidth area. Finally, the Ramulator provides a fixed 25 ns latency in the whole bandwidth area and for all memory traffic configurations. Also, similar to the fixed-latency model, the simulated bandwidth is unrealistic, exceeding by 1.8$\times$ the maximum theoretical one.
Our evaluation of the memory models and detailed hardware simulators detected major differences w.r.t. the actual memory systems performance. DRAMsim3 and Ramulator, which rely on detailed memory models, are considered de facto standard for the memory system simulation. Also, both of them are evaluated against the manufacturer’s Verilog implementations and they show no violation of the JEDEC timings [55]. However, as our results demonstrate, this does not guarantee that the simulators properly model the full system performance.
II-E3 RTL simulators: OpenPiton and Metro-MPI
In our experiments, the OpenPiton framework is configured to generate 64-core Ariane architecture which includes 16 KB of 4-way L1 instruction and data cache, and 4 MB of 4-way shared L2 cache. The main memory is originally modeled with a single-cycle latency, and it is recently extended with a fixed-latency model [74]. We set this latency to 170 ns. Our Mess measurements, see Figure 6, confirm that both models deliver the expected load-to-use latency. Also, as expected, we see no difference between read and write memory traffic, leading to a perfect overlap of the curves. The only difference is in the maximum observed memory bandwidth. For a single-cycle memory latency, 100%-read memory traffic achieves 32 GB/s, limited by the memory concurrency of the 64 in-order Ariane cores. Memory writes do not stall the cores, so the achieved memory bandwidth increases with the write memory traffic ratio. Still, a small 2-entry miss status holding registers (MSHR) limits the memory bandwidth to 47 GB/s for 50%-read/50%-write traffic. We detect the same trend for the fixed memory model.
The Mess evaluation of the OpenPiton Metro-MPI resulted in an unexpected discovery: in some experiments we detected significantly higher memory write traffic than anticipated. By analysis of the system behavior for various Mess configurations, we connected the extra memory traffic to the unnecessary eviction of the data from the last-level cache. Instead of evicting only the dirty cache lines, the system was evicting all of them. The source of the error is the coherency protocol generated by the OpenPiton framework. The error is reported to the OpenPiton developers and they confirmed its existence. Although not part of the original plan, our exploration discovered that holistic and detailed Mess performance characterization can be also used to uncover CPU design errors.
(a) Single cycle model
(b) Fixed-latency model
Figure 6: Memory system performance of the OpenPiton Metro-MPI simulator.
III Mess simulator

### v4

@@ III Performance characterization: Actual systems
All platforms except AMD Zen2 CPU and NVIDIA H100 show a similar saturated bandwidth range, between approximately 70% and 90% of the maximum theoretical bandwidth. The maximum achieved bandwidth cannot reach the theoretical one because part of it is “lost” due to factors such as DRAM refresh cycles blocking the entire chip, page misses causing precharge and activate cycles, and timing restrictions at the bank, rank, and channel levels [66]. The efficiency of the memory bandwidth utilization also depends on the memory controller design. As expected (see Section II-C), the best utilization is achieved for 100%-read memory traffic, and it reduces with the increment of the memory writes. AMD Zen2 is an exception in two ways. First, its saturated bandwidth range is significantly lower, 57–71% of the maximum theoretical one. Second, it does not follow the expected impact of the write traffic on the bandwidth utilization. The traffic with the maximum rate of memory writes shows a very good performance, very close to the 100%-read traffic, while the main drop is detected for a mixed, e.g. 60%-read/40%-write, traffic.
@@
In some AMD Zen 2, Intel Skylake and Intel Cascade Lake bandwidth–latency curves, increase in the Mess memory access rate, after some point, leads to decline in the measured bandwidth. [^2] These findings are consistent with two recent studies which report that running high-bandwidth benchmarks on all CPU cores may lead to lower memory bandwidths w.r.t. the experiment in which some cores are not used [6, 67]. In Amazon Graviton 3, Intel Sapphire Rapids and NVIDIA H100 this behavior is frequent for the memory traffic with high percent of writes. We explored these findings in the Cascade Lake servers in which we had access to the row-buffer hardware counters. In the experiments that experienced the bandwidth decline, we detect a significant increase in row-buffer misses. In case of a row-buffer miss, the current content of the row is stored in the memory array and the correct row is loaded into the row-buffer. These additional operations increase the memory access time and reduce the effective device bandwidth. To confirm these findings, we increased the memory system pressure by removing four out of six DIMMs in each socket. In these experiments, we detected a large “wave-form” segments in all bandwidth–latency curves. And indeed, the measurement confirmed that the measured bandwidth drop is highly correlated with the row-buffer miss rate.
@@
The results also show the difference between the STREAM and Mess measurements. STREAM benchmark reports four data-points: the maximum sustained bandwidth of Copy, Scale, Add and Triad kernels. Mess provides a wide range of measurements for different memory traffic intensity and read/write ratios. STREAM measures the application-level data traffic estimated based on the application execution time, size of the data structures, and number of load and store instruction in each STREAM kernel. Mess measures the architecture-level memory bandwidth measured with hardware counters. This includes all the application data traffic, but also all the microarchitecture data transfers. For example, STREAM bandwidth calculation assumes one memory read for each load instruction and one memory write for each store. In state-of-the-art HPC servers with write-allocate cache policy, each store instruction does not correspond to a single memory write, but to one read and one write (see Section II-A). For this reason, Mess benchmark reports higher maximum measured bandwidths. This difference is clearly visible in IBM Power 9, A64FX and all Intel platforms. In Graviton 3 and NVIDIA H100, the results reported by STREAM are very close to the maximum Mess measurements for the corresponding read/write ratio. This would correspond to a architecture with a write-through cache policy. The second cause of the difference between the maximum Mess and STREAM bandwidths is the memory traffic composition. The Mess benchmark achieves the maximum bandwidth for a 100%-read traffic. The write memory traffic, present in all STREAM kernels, adds timing constrains and reaches sooner the bandwidth saturation point. Our results show that different platforms show distinct relation between the application-level STREAM and architecture-level Mess measurements. Therefore, we believe that the memory bandwidth analysis should include both approaches. As a part of future work, we plan to compare the STREAM and Mess measurements in high-end architectures that dynamically adjust the level of write-allocate traffic [68, 69].
IV Performance characterization: Memory simulators
@@
Mess benchmark can also be used to characterize memory simulators and compare them with the actual systems they intent to model. We illustrate this Mess capability with the gem5, ZSim, and OpenPiton Metro-MPI simulators with different internal memory models and widely-used external memory simulators, DRAMsim3, Ramulator and Ramulator 2.
IV-A gem5
@@
The gem5 [16] is a cycle-accurate full-system simulator. In our experiments, the simulator is configured to model the Graviton 3 server with 64 Neoverse N1 cores [61]. The cache hierarchy includes 64 KB of 4-way L1 instruction and data cache, 1 MB of 8-way private L2 cache and 64 MB of 16-way shared L3. The main memory system has eight DDR5-4800 memory channels. Figure 4 compares Mess bandwidth–latency curves of the actual server with gem5 simple memory model, more complex gem5 internal DDR model and gem5 connected to Ramulator 2. To maintain a reasonable simulation time, we model each system with a family of six curves, from 50% to 100% read memory traffic with a 10% step.
@@
Practically in the whole bandwidth range, the gem5 simple memory model delivers a fixed latency of 4–49 ns. The latency increases only when the bandwidth asymptotically approaches its theoretical maximum. Contrary to the Graviton 3 server, the highest latencies are measured for a 100%-read traffic, and the latency drops with the percent of memory writes. Also, unlike in the actual platform, for some memory traffic, increasing the bandwidth reduces the memory access latency. For example, the 50%-read/50%-write traffic reaches the lowest simulated latency of only 4 ns at the 200 GB/s bandwidth. The same traffic in the actual system has the memory access latency of 261 ns.
@@
The more detailed internal DDR model shows small improvements over the simple memory model, but still poorly resembles the actual system performance. The simulated latencies are unrealistically low, most of them in the range of 14–100 ns. Similarly to the gem5 simple memory model, the latencies drops with the percent of memory writes. For all the curves except 100%-read, the saturated bandwidth is significantly lower from the one measured on the actual system. Again, the error increases with the percent of memory writes.
As the internal gem5 memory models, gem5+Ramulator 2 simulates unrealistically low memory latencies and the error increases with the ratio of memory writes. In addition to this, the curves experience a sharp, nearly vertical rise between 100 GB/s and 130 GB/s, which is less than a half of the actual measured bandwidth. Surprisingly, the most complex and trusted memory model shows the highest simulation error. Some sources of this error will be analyzed in Section IV-D.
@@
(d) gem5+Ramulator2
Figure 4: Memory performance: Amazon Graviton3 server vs. gem5 memory models.
IV-B ZSim
We select ZSim [15] as a representative of event-based hardware simulators. We use publicly-available ZSim modeling 24-core Intel Skylake processor connected to six DDR4-2666 channels [70]. The cache hierarchy of the modeled CPU includes 64 KB of 8-way L1 instruction and data cache, 1 MB of 16-way private L2 cache and 33 MB of 11-way shared L3. The simulator is extensively evaluated against the actual hardware platform [71]. The ZSim comprises three internal memory models: fixed-latency, M/D/1 queue model and the internal DDR model. Also it is already connected to Ramulator [23] and DRAMsim3 [22]. To avoid any simulator integration error, we use the ZSim+DRAMSim3 released by the University of Maryland (DRAMSim3 developers) [72], and ZSim+Ramulator from The SAFARI Research Group at the ETH University (Ramulator developers) [73].
Figure 5 compares the Mess bandwidth–latency curves of the actual server with all five ZSim memory simulation approaches. As expected, the fixed-latency memory model provides a constant latency in the whole bandwidth domain. Given that this latency is configured by a user, it can be set to match the unloaded memory latency in the actual system. On the down side, the memory bandwidth provided by this model is unrealistic: the maximum simulated bandwidth is 342 GB/s, which exceeds the maximum theoretical one by 2.7$\times$. The M/D/1 queues correctly model the memory system behavior in the linear part of the curves. The modeling of the system saturation is less accurate. The queue model does show some difference between read and write memory traffic, but the reported performance does not correspond to the actual system trend in which increasing the write traffic lowers the performance. The internal DDR model correctly emulates the linear and saturated segments of the curves and the impact of the memory writes. However, the simulator underestimates the saturated bandwidth area to 69–93 GB/s, significantly below the 92–116 GB/s measured in the actual system. Also, the simulator excessively penalizes the memory writes which is seen as a wider spread of the curves with a higher write memory traffic. Finally, we detect some unrealistic memory-latency peaks in the low-bandwidth 1–4 GB/s curve segments. The DRAMsim3 shows a similar trend as the M/D/1 queue model in the linear segments of the memory curves, with some latency error, 52–63 ns in the DRAMsim3 versus 89–109 ns in the actual system. The simulator does not model the saturated bandwidth area. Finally, the Ramulator provides a fixed 25 ns latency in the whole bandwidth area and for all memory traffic configurations. Also, similar to the fixed-latency model, the simulated bandwidth is unrealistic, exceeding by 1.8$\times$ the maximum theoretical one.
Our evaluation of the memory models and detailed hardware simulators detected major discrepancies w.r.t. the actual memory systems performance. DRAMsim3, Ramulator and Ramulator 2 are considered de facto standard for the memory system simulation. All of them are evaluated against the manufacturer’s Verilog model and they show no violation of the JEDEC timings [74, 57]. DRAMsim3 is evaluated for DDR3 and DDR4 [22], Ramulator for DDR3 [23], and Ramulator 2 for DDR4 [24]. However, as our results demonstrate, this does not guarantee that the simulators properly model the memory system performance. In Section IV-D, we will use Mess benchmark to analyze some causes of these discrepancies.
IV-C RTL simulators: OpenPiton and Metro-MPI
In our experiments, the OpenPiton framework is configured to generate 64-core Ariane architecture which includes 16 KB of 4-way L1 instruction and data cache, and 4 MB of 4-way shared L2 cache. The main memory is originally modeled with a single-cycle latency, and it is recently extended with a fixed-latency model [77]. Our Mess measurements confirm that both models deliver the expected load-to-use latency. Also, as expected, we see no difference between read and write memory traffic, leading to a perfect overlap of the curves. The only difference is in the maximum observed memory bandwidth. For a single-cycle memory latency, 100%-read memory traffic achieves 32 GB/s, limited by the memory concurrency of the 64 in-order Ariane cores. Memory writes do not stall the cores, so the achieved memory bandwidth increases with the write memory traffic ratio. Still, a small 2-entry miss status holding registers (MSHR) limits the memory bandwidth to 47 GB/s for 50%-read/50%-write traffic. We detect the same trend for the fixed memory model.
The Mess evaluation of the OpenPiton Metro-MPI resulted in an unexpected discovery: in some experiments we detected significantly higher memory write traffic than anticipated. By analyzing the system behavior for various Mess configurations, we connected the extra memory traffic to the unnecessary eviction of the data from the last-level cache. Instead of evicting only the dirty cache lines, the system was evicting all of them. The source of the error is the coherency protocol generated by the OpenPiton framework. The error was reported to the OpenPiton developers and they confirmed its existence.
@@ IV-C RTL simulators: OpenPiton and Metro-MPI
Figure 5: Memory performance: Intel Skylake server vs. ZSim memory models.
IV-D Sources of memory simulation errors
To exclude any simulation error caused by the CPU simulators or their memory interfaces, we perform a trace-driven DRAMsim3, Ramulator and Ramulator 2 simulation. The detailed Mess memory traces are collected from its ZSim simulation, and they include the addresses of all memory read and write operations. To account for the timings of non-memory operations, DRAMsim3 traces contain simulation cycles in which the memory requests reach the memory controller, The Ramulator and Ramulator 2 traces include the number of non-memory instructions between the consecutive memory operations. The Mess traces may contain some timings errors w.r.t. actual execution, so the DRAMsim3, Ramulator, and Ramulator 2 simulations may not match the exact bandwidth–latency point. Still, the correct memory simulation should provide data-points that are located on the actual bandwidth–latency curves. Figure 6 shows the trace-driven Mess evaluation of DRAMsim3, Ramulator and Ramulator 2. The charts report the round-trip memory access latency from the memory controller, so it is expected that the simulated curves are somewhat below the actual load-to-use measurements.
The trace-driven Ramulator 2 show the same bandwidth–latency trends as the gem5+Ramulator 2 simulations (see Figure 4(d)) . The simulated memory latency is unrealistically low and the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system. This indicates that the main source of the large simulation error is indeed Ramulator 2.
The conclusions are somewhat different for DRAMsim3 and Ramulator as their trace-driven memory bandwidth–latency curves show better general trends than the corresponding Zsim-driven simulations. This indicates that a part of the ZSim+DRAsim3 and ZSim+Ramulator simulation errors reported in Figure 5 is caused by the simulators’ interfaces. Our finding is aligned with the previous studies that report issues in the integration of the event-based CPU simulators with cycle-accurate memory models [49, 80, 81].
However, trace-driven DRAMsim3 and Ramulator also show important discrepancies w.r.t. actual bandwidth–latency curves. DRAMsim3 simulated latency starts at 68 ns. Apart from the peak at 5 GB/s, the latency increases linearly with the bandwidth. The curves for different read/write ratios are spread and intertwined in whole bandwidth range. Below 70 GB/s the curves with the highest write traffic ratio have the lowest latency. We detect no bandwidth saturation, and all the curves linearly reach the maximum bandwidth of 113 GB/s. Ramulator shows a better general trend: roughly-constant latencies below 40 GB/s, a light latency increase until approximately 85 GB/s, and a higher inclination in the final segments of the curves. Still, the latency is unrealistically low, starting at only 25 ns (100%-read traffic, 20–40 GB/s), different read/write curves are spread in all bandwidths, and the saturated behavior differs significantly from the actual one.
(a) Ramulator2: 8$\times$×DDR5-4800
(b) DRAMsim3: 6$\times$×DDR4-2666
(c) Ramulator: 6$\times$×DDR4-2666
Figure 6: Memory performance: Trace-driven cycle-accurate simulators.
To understand better the underlying causes of the trace-driven DRAMsim3 and Ramulator simulation errors, we compare their row-buffer hit, empty and miss statistics with to the measurements from the actual Intel platform. We could not perform the same analysis for the Ramulator 2 because its baseline architecture, Amazon Graviton3 with 8$\times$×DDR5-4800, does not support the row-buffer measurements. Figure 7 shows a subset of the results for the 100%-read and 50%-read/50%-write memory traffic, which is sufficient to show a general trend. For a 100%-read traffic and low memory bandwidth utilization, the actual system has 84% row-buffer hits, 13% empty buffers and 3% misses. As expected, higher memory bandwidth utilization decreases the hit ratio, and increases the empty pages and misses. Also, as we increase the write traffic, the row-buffer utilization degrades, compare Figure 7(a) and 7(b).
(a) 100%-read
(b) 50%-read/50%-write
Figure 7: Row buffer statistics: Actual hardware vs. DRAMsim3 vs. Ramulator.
DRAMsim3 shows a very different behavior. In most of the experiments, we measure 84–93% row-buffer hit rate with 7–16% of the misses. We detect the highest hit-rates for the dominantly-read and dominantly-write traffic, while intermediate read/write ratios have lower values. These hit-rates match the vertical spread we see in the DRAMsim3 bandwidth-latency curves for low and medium bandwidths (80 GB/s). The curves with high hit-rates have the lowest latencies. For the curves with lower hit-rates, the latency increase by up to 20 ns. In 2 GB/s DRAMsim3 experiments some of the read/write ratios have a surprisingly low row-buffer hit-rates ($<35\%$), perfectly matching the Figure 6(b) memory latency peak at this bandwidth data-point. The Ramulator row-buffer statistics resemble better the actual measurements, see Figure 7. Still we detect some discrepancies that, interestingly, have similar trends as the DRAMsim3 results. Again, we detect the highest hit-rates for the dominantly-read and dominantly-write traffic, while intermediate read/write ratios have lower values. For $>40\%$ write traffic, Ramulator hit-rates greatly exceed the actual ones in the whole bandwidth range (Figure 7(b)). As in DRAMsim3, these hit-rates closely match the vertical spread of the Ramulator latency simulations.
Comparison of the actual platform row-buffer statistics and the latency measurements resulted in an interesting finding. For low and moderate bandwidth utilization ($<$70 GB/s), the write traffic increase leads to a notably worse row-buffer utilization, but this does not translate into higher memory access latencies. It seems that the actual system is capable to mask the row-buffer contentions and delays. We do not see this behavior in the DRAMsim3 and Ramulator simulations. Overall, our analysis of the row-buffer statistics and its correlation with the bandwidth–latency curves provides some first steps in the analysis of the memory simulation errors. We believe that our work and publicly-released Mess benchmark will motivate the community to continue this exploration.
The memory access pattern has a significant impact on the row-buffer utilization and the overall memory system performance. The Mess benchmark concurrently executes random-access pointer-chase and the multi-process memory traffic generator. Each process traverses two arrays, one with load and one with store operations. Each array is accessed sequentially, but, the overall memory access pattern is complex due to the concurrent traversal of tens of distinct arrays located in main memory. Mess traffic generator can be easily extended to cover different array access patterns. Some of these patterns are strided access, e.g. targeting a new row-buffer in each operation, or the random access, e.g. the RandomAccess test in HPC Challenge benchmark developed to measure Giga Updates Per Second (GUPS) in a system [82].
V Mess simulator